How long will it take for a 40.0 gram sample of I-131 (half life = 8.040 days) to decay to 1/100 its original mass?
1
Expert's answer
2014-10-30T15:22:35-0400
1/100= 0.01. Suppose that n – is number of half lifes, then (1/2)n = 0.01 then n = 6.64. So, the time is 6.64*8.040 = 53.4days.